OpenStudy (anonymous):

Evaluate: log4(8)+ log4(16) – log4(2)

OpenStudy (solomonzelman):

\[\log_48+\log_416-\log_42=\log(8 \times 16 \div 2)=\log_464=\log_44^3=3\log_44=2 \times 1 =3\]

OpenStudy (solomonzelman):

Where it say 2 *1 it's 3 *1

OpenStudy (solomonzelman):

it's 3.